how do you simplify the expression 40, using positive exponents?
Answer by tommyt3rd(5050)

when you say simplify do you mean factor?
if yes then 40=2*20=2*2*10=2^3*5^1
otherwise...
what needs to be done is not clear to me from the given information please repost with a clear and complete question :)